Sol Summary – February 5th

Summary title : Mangalyatri, Crew 292 Begins Simulation After Training.
Author’s name: Annalea Beattie
Mission Status: Nominal
Sol Activity Summary: Despite delays and snowy stormy weather Mars Society Australia Crew 292 Mangalyatri (Mars Explorers) all finally arrived at the Mars Desert Research Station, joining crew member Clare Fletcher who was also a member of MSA Crew 291. This morning we trained for simulation and learnt about all the systems of the Mars hab. We spent time in the engineering bay (RAM – the repair and maintenance module), in the fruiting and plentiful Green Hab and in the extraordinary Science Dome. Later, outside, MDRS Director Sergii Lakymov taught us how to charge and drive the electric rovers and we all went for a practice drive. For our crew geologist Bharti Sharma, it was her first experience ever behind the wheel of a vehicle. In the afternoon we were trained in comms with radios and headsets and then we tried on space suits. This was great and congratulations, North California Mars Society – the suits you have made for the station are excellent, so comfortable and easy to wear. At 3.30 pm everyone sat down for a cup of tea and a quick planning meeting. We discussed tomorrow’s two training EVA’s, and our dinner roster for this evening. We examined our water use and thought about how we can conserve our resources. We scheduled a research discussion meeting for tomorrow and decided that we would begin simulation at six pm today. With two hours to go of fresh air outside, we all went for a walk up to Hab Ridge. The Henry Mountains are snow-covered and beautiful and we wandered along the top of the Dakota sandstone. We came back to the hab via spectacular North Ridge with its deco colours – creamy pinks, chocolate browns and ochres, rusts and black. The desert was quiet and still, the ground soft from the rain and we walked home on pink mud like Mars. We came inside, shut the doors and began simulation.
Look Ahead Plan: Two training EVA’s and a research project meeting to roster in field work.
Anomolies in work: Nothing to report
Weather: Light cloud cover today with mild wind, NNE 8-17 mph. Daily temperature 51-54F.
Crew Physical Status: Nominal
Reports to be filed: Sol Summary, Journalist Report, Greenhab Report, Operations Report, EVA report, EVA Request(s)
Support requested: Nothing to request.

Copyright © The Mars Society. All rights reserved. | Main Site